# Enhancing Vision-Language Compositional Understanding with Multimodal Synthetic Data

### v1

@@ Abstract
Despite impressive advancements in various multimodal tasks, vision-language models (VLMs) still struggle with compositional understanding due to limited exposure to training samples that contain subtle variations within paired examples. With advances in multimodal generative models, a natural solution is to generate synthetic samples with subtle variations for training VLMs. However, generating and training on synthetic samples with subtle variations presents two challenges: difficulty in accurately creating precise variations and inconsistency in cross-modal alignment quality. To address these challenges, we propose SVD-GT (Subtle Variation Data Generation and Training), which integrates image feature injection into a text-to-image generative model to enhance the quality of synthetic variations and employs an adaptive margin loss to differentiate samples using adaptive margins, which help filter out potentially incorrect synthetic samples and focus the learning on informative hard samples. Evaluations on four compositional understanding benchmarks demonstrate that SVD-GT significantly improves the compositionality of VLMs, boosting the average accuracy of CLIP by over 8% across all benchmarks and outperforming state-of-the-art methods by 2% on three benchmarks.
@@ 1 Introduction
Recent advancements in large-scale pretrained Vision-Language Models (VLMs) [71, 40, 99, 81, 1, 49, 50, 114] have significantly enhanced performance across a wide range of multimodal tasks [113, 112, 96, 14]. However, VLMs still face limitation in accurately interpreting compositional relationships between objects and attributes, as demonstrated by various evaluations [85, 110, 103, 58, 32]. This limitation primarily stems from the absence of subtle variations within paired samples in the training data [41]. As a result, VLMs often depend on shortcut features [24] for image-caption alignment, rather than genuinely learning nuanced distinctions. While collecting training samples with subtle variations could enhance compositionality, this approach is time-consuming and labor-intensive, making it impractical at scale.
@@
However, generating and training on synthetic data presents two key challenges. The first is the difficulty of efficiently and accurately creating precise variations in synthetic images. To generate large-scale training data with precise variations, the image generation process must meet three criteria: efficiency in producing large quantities of images, text alignment between generated images and their corresponding captions, and image fidelity in ensuring the generated images closely match their real counterparts. Nevertheless, current image generation methods struggle to meet all of these criteria simultaneously. Per-sample optimization methods [43, 91, 109, 79] lack efficiency, while zero-shot image editing methods [59, 8, 23] and text-to-image (T2I) models [72, 75] often fail to achieve proper text alignment and image fidelity, respectively, as illustrated in Figure 1 (a). The inaccurate synthetic variations produced by these models can mislead the learning of VLMs. To alleviate this issue, we propose injecting real image features into the text prompt features of a fast T2I model, which excels in efficiency and text alignment, to enhance image fidelity. By combining this approach with AdaIN [36], the image generation process largely meets all three criteria, enabling the generation of subtle variations in images.
Another challenge is the inconsistency in cross-modal alignment quality. Due to the limitations of generative models in fully adhering to prompts [39, 69] or generating precise content [35, 4], synthetic samples often exhibit varying degrees of cross-modal alignment. The similarity between positive and negative pairs varies, with some being hard pairs, others easy, and some even representing incorrect generations, as shown in Figure 1 (b). Uniformly treating all positive and negative samples fails to account for the variability in alignment quality. To leverage the varying degrees of cross-modal alignment, we propose an adaptive margin loss that differentiates between positive, hard negative, and easy negative samples, with adaptive margins. The adaptive margins help filter out potentially incorrect generations and focus the learning on informative hard samples.
@@
We propose SVD-GT, which integrates image feature injection into a fast T2I model to improve the quality of synthetic variations in images and employs an adaptive margin loss to leverage the varying alignment quality in synthetic samples for training VLMs. Evaluations on four compositional understanding benchmarks demonstrate that SVD-GT significantly enhances VLM compositionality, improving CLIP by over 5% on VL-CheckList [110] and 7% on SugarCrepe [32], while surpassing state-of-the-art methods by 1% and 2% on the two benchmarks. The main contributions of this paper are as follows: (1) We propose image feature injection to enhance the quality of synthetic variations in images, which provide valuable training data to improve the compositionality of VLMs. (2) We introduce an adaptive margin loss to leverage varying levels of cross-modal alignment in synthetic samples to effectively differentiate positive and negative samples. (3) Experimental results confirm that SVD-GT significantly improves the compositional understanding capabilities of VLMs.
@@ 2.1 Limitations in Compositionality of VLMs
While VLMs excel in many multi-modal tasks [71, 40, 99, 81, 49, 50, 114], they still struggle with compositional understanding—the ability to interpret novel combinations of known visual and textual components. Benchmarks like What’sUp [41] reveal difficulties in understanding spatial relationships, SPEC [66] highlights issues with object size, position, and count, and ARO [103] uncovers limitations in understanding attributes, relations, and word order. Winoground [85], SNARE [93], and VL-CheckList [110] also expose these shortcomings. SugarCrepe [32] addresses hackable biases in prior benchmarks, where text-only models achieve artificially high performance, by introducing fluent and meaningful hard negatives. Their findings suggest that previous benchmarks overestimated compositional understanding. Building on this, SugarCrepe++ [20] further introduces semantically equivalent but lexically varied captions as hard positives and shows the difficulties of VLMs in distinguishing between lexical and semantic variations.
@@ 2.2 Improving Compositionality of VLMs
Prior approaches to improving the compositionality of VLMs can be broadly classified into the following categories: (1) Leveraging detailed image captions: Detailed captions from dense captioning models [18, 51], simulation platforms [11], and video annotations [44] are collected to train VLMs. However, these samples often lack pairs with subtle variations, limiting their contribution to compositionality. (2) Distilling from pretrained models: SDS-CLIP [5], SF-CLIP [77] and IL-CLIP [111] distill knowledge from pretrained image generation models and visual-language foundation models. IL-CLIP [111] refines representations through iterative learning with pretrained vision and language agents. However, pretrained models also face limitations in compositionality [34, 87]. (3) Incorporating structural knowledge: MosaiCLIP [82] and Structure-CLIP [37] incorporate scene graph knowledge in text features. CLIP-SGVL [28] and 3VL [100] train VLMs to predict scene graphs. [62] utilizes scene graphs as prompts to elicit compositional knowledge from VLMs without further training. However, these methods rely on models trained with expensive dense structure annotations (e.g., scene graphs). (4) Utilizing synthetic negative samples: Rule-based tools [30] or large language models (LLMs) [16, 70] are used to generate negative captions by editing real captions [103, 19, 12, 83, 80, 64, 106]. Image generation models [72] are also used to create or edit images for training [76, 67, 105, 48]. Despite their utility, efficiently generating large amount of images with precise variations remains challenging due to the inherent limitations of generative models. (5) Applying fine-grained alignment constraints: MCD [45] enforces multi-scale alignment across images with varying augmentations and the corresponding text captions. SPARC [7] learns local alignment by associating each text token with a group of local image patches. CE-CLIP [106] applies intra-modal contrastive loss and cross-modal ranking loss to improve alignment. However, uniform supervision applied to synthetic samples with varying alignment quality limits their effectiveness. In this paper, we propose SVD-GT to address two key challenges in learning from synthetic data: the difficulty in accurately creating precise variations and the inconsistency in cross-modal alignment quality in synthetic data.
@@ 2.3 Training with Synthetic Data
The use of synthetic data for training machine learning models is common across various fields [63, 73, 74, 47, 97, 27, 60, 15, 90]. Synthetic data generated through simulations and graphics engines supports a wide range of tasks, including optical flow estimation [17], domain adaptation [68] and action recognition [92]. However, these synthetic datasets often diverge significantly from real-world data. Recent advances in generative models [9, 107, 88, 42, 13, 72, 75, 102, 57] have made it possible to synthesize data that more closely resembles real-world scenarios. In language tasks, synthetic textual data are widely used [22, 31, 60, 61, 94, 95, 115, 84, 98, 56]. In computer vision, synthetic data assist in tasks such as image classification [6, 26, 2, 78], object detection [29, 89, 108], and unsupervised visual representation learning [38, 53, 3, 86]. In this paper, we generate multimodal samples with subtle variations to enhance the compositionality of VLMs.
Figure 2: An overview of SVD-GT. (a) Starting with a real image-caption pair, we generate synthetic positive and negative pairs with subtle variations using an LLM and a fast T2I model. To improve the quality of subtle variations in synthetic images, we introduce image feature injection to reduce unintended variations from a standard T2I model (see Sec. 3.2). (b) We train the VLM using both real and synthetic samples. In addition to a sigmoid loss for distinguishing positive and negative pairs, we apply an adaptive margin loss that leverages varying alignment levels across training samples to learn informative nuanced distinctions (see Sec. 3.3).
3 SVD-GT
To improve the compositional understanding abilities of VLMs, we propose SVD-GT, which generates multimodal samples with subtle variations from real samples and trains VLMs to learn nuanced differences through synthetic data. In the generation phase, SVD-GT creates positive and negative captions with slight variations from real captions using an LLM, then generates images based on the real image and these modified captions through a fast T2I model. To enhance the quality of subtle variations in synthetic images, we introduce image feature injection, which integrates real image features into the text prompt features of the T2I model to improve fidelity to the real image, as detailed in Sec. 3.2. In the training phase, SVD-GT employs an adaptive margin loss that leverages varying levels of multimodal alignment in the synthetic samples to effectively learn informative nuanced distinctions, as described in Sec. 3.3. The framework of SVD-GT is shown in Figure 2.
@@ 3.1 Generating Negative and Positive Captions
Captions with subtle variations are crucial for learning compositional knowledge, as shown by previous work that generates captions by randomly swapping or replacing nouns and adjectives [103, 106]. However, manually designed generation rules often introduce nonsensical or grammatically incorrect artifacts, creating shortcut features [24] that obstruct true compositional understanding [32]. To address this issue, we use an LLM to generate natural synthetic captions. To further mitigate the impact of generative artifacts, we generate both negative and positive captions, ensuring that VLMs cannot easily differentiate negative captions from positive ones based solely on artifacts. We denote the $i^{th}$ real image-caption pair in the training set as $(\bm{I}_{i}^{r},\bm{T}_{i}^{r})$. Given a real caption $\bm{T}_{i}^{r}$, we prompt the LLM to generate a synthetic negative caption $\bm{T}_{i}^{sn}$ and a synthetic positive caption $\bm{T}_{i}^{sp}$, as specified by the prompts in Figure A1 in Appendix.
@@ 3.2 Generating Images via Image Feature Injection
Given a real image $\bm{I}_{i}^{r}$ and the corresponding synthetic caption $\bm{T}_{i}^{s}$ ($s\in\{sn,sp\}$), we extract the image embedding $\bm{f}_{i}^{r}$ (the [CLS] token) and the text embedding $\bm{e}_{i}^{s}=\langle\bm{e}_{i,j}^{s}\rangle_{j=1}^{L}$, where $L$ is the maximum text length, using two aligned feature encoders (e.g., CLIP image and text encoders). Let the index of the [EOS] token in $\bm{e}_{i}^{s}$ as $k_{i}$. We replace the padding embeddings with the real image embedding to produce a new embedding:
$\hat{\bm{e}}_{i}^{s}=\left\{\begin{array}[]{ll}\bm{e}_{i,j}^{s},&j\leq k_{i}\\ \bm{f}_{i}^{r},&j>k_{i}\end{array}\right.$ (1)
The embedding $\hat{\bm{e}}_{i}^{s}$ is then used as input to the T2I model, which generates a synthetic image $\tilde{\bm{I}}_{i}^{s}$. This process, depicted in Figure 2, reduces unintended variations in synthetic images generated by the standard T2I model and enhances image fidelity to real images, thereby improving the quality of synthetic subtle variations.
Style Transfer. To further reduce the domain gap between synthetic and real images, we use AdaIN [36] to transfer the style of the synthetic image to that of the real image. A pretrained encoder-decoder with AdaIN [36] is employed to align the channel-wise features of $\tilde{\bm{I}}_{i}^{s}$ with those of $\bm{I}_{i}^{r}$, resulting in $\bm{I}_{i}^{s}$, which is then used for model training.
@@ 3.3 Training with Real and Synthetic Samples
After generating captions and images, each real image-caption pair $(\bm{I}_{i}^{r},\bm{T}_{i}^{r})$ is extended to $(\bm{I}_{i}^{r},\bm{T}_{i}^{r},\bm{I}_{i}^{sn},\bm{T}_{i}^{sn},\bm{I}_{i}^{sp},\bm{T}_{i}^{sp})$ by adding one synthetic negative and one synthetic positive pair. We then train the VLM using these extended samples. Given a batch of $n$ sample groups, $\{(\bm{I}_{i}^{r},\bm{T}_{i}^{r},\bm{I}_{i}^{sn},\bm{T}_{i}^{sn},\bm{I}_{i}^{sp},\bm{T}_{i}^{sp})\}_{i=1}^{n}$, we organize the images by concatenating all real, synthetic negative, and synthetic positive images into an image batch, $\bm{I}^{B}=\langle\bm{I}_{1}^{r},\cdots,\bm{I}_{n}^{r},\bm{I}_{1}^{sn},\cdots,\bm{I}_{n}^{sn},\bm{I}_{1}^{sp},\cdots,\bm{I}_{n}^{sp}\rangle\in\mathbb{R}^{3n\times h\times w\times c}$, where $h$, $w$, and $c$ represent the height, width, and number of channels of the images, respectively. Similarly, we organize all captions in the same order into a list, $\bm{T}^{B}$, of length $3n$.
We calculate the similarity between each image-caption pair in the batch. The similarity between the $i^{\text{th}}$ image $\bm{I}_{i}^{B}$ and the $j^{\text{th}}$ caption $\bm{T}_{j}^{B}$ is given by $S_{i,j}=c(\mathcal{E}_{I}(\bm{I}_{i}^{B}),\mathcal{E}_{T}(\bm{T}_{j}^{B}))$, where $c(\cdot,\cdot)$ denotes cosine similarity, and $\mathcal{E}_{I}$ and $\mathcal{E}_{T}$ represent the image and text encoders of the VLM, respectively. We then define the ground-truth assignment matrix $\bm{M}\in\{0,1\}^{3n\times 3n}$ as follows:
$M_{i,j}=\left\{\begin{array}[]{ll}1,&(j=i)\lor(j=i+2n)\lor(j=i-2n),\\ -1,&\text{otherwise}.\end{array}\right.$ (2)
The matrix $\bm{M}$ encodes the correct alignments between images and captions within the batch, where an entry $M_{i,j}=1$ represents a positive match, and $M_{i,j}=-1$ denotes a non-matching pair. To account for cases where an image or caption may have multiple positive associations, we apply a sigmoid-based contrastive loss [104, 10] to encourage higher similarity for positive pairs and lower similarity for negative pairs. The contrastive loss is computed as follows:
@@
Adaptive Margin Loss. Synthetic samples often lack consistent cross-modal alignment due to the difficulties in precisely controlling generative models. For instance, in Figure 1 (b), the three examples display noticeably different similarity levels between positive pairs and negative pairs. Applying uniform positive or negative labels to all samples in Eq. (3) fails to account for this variation in alignment quality, potentially misleading the VLM with wrongly generated samples or failing to focus on informative hard samples. To leverage the varying cross-modal alignment levels during training, we propose an adaptive margin loss that differentiates positive and negative samples with adaptive margins, which help filter out potential incorrect samples and prioritizes learning from hard samples.
For the $i^{\text{th}}$ image in a batch ($i={1,\dots,3n}$), we define four sets that represent different levels of alignment based on the sample arrangement within the batch. These sets are: a positive set $\mathbb{P}$, a hard negative set $\mathbb{N}_{h}$, an easy negative set $\mathbb{N}_{e}$, and a real negative set $\mathbb{N}_{r}$. The $j^{\text{th}}$ caption from each set corresponds to the positive caption, hard negative caption, easy negative caption, and real negative caption, respectively. The definitions of these sets are as follows:
[…]
We then define the adaptive margin loss for the $i^{th}$ image as follows:
$\begin{split}L_{mar,i}^{I}&=\frac{1}{\lvert\mathbb{P}\rvert\cdot\lvert\mathbb{N}_{n}\rvert}\sum_{j_{1}\in\mathbb{P},j_{2}\in\mathbb{N}_{n}}\text{max}(0,m+S_{i,j_{2}}-S_{i,j_{1}})\\ &+\frac{1}{\lvert\mathbb{N}_{h}\rvert\cdot\lvert\mathbb{N}_{e}\rvert}\sum_{j_{1}\in\mathbb{N}_{h},j_{2}\in\mathbb{N}_{e}}\text{max}(0,m+S_{i,j_{2}}-S_{i,j_{1}})\\ &+\frac{\alpha}{\lvert\mathbb{P}\rvert\cdot\lvert\mathbb{N}_{r}\rvert}\sum_{j_{1}\in\mathbb{P},j_{2}\in\mathbb{N}_{r}}\text{max}(0,m+S_{i,j_{2}}-S_{i,j_{1}}).\end{split}$ (5)
This loss encourages positive pairs to have higher similarity scores than hard negative pairs, and hard negative pairs to have higher similarity than easy negative pairs. Additionally, a weight $\alpha>1$ is applied to comparisons between positive pairs and real negative pairs to emphasize these comparisons, as real samples are generally correct, making comparisons involving them more reliable.
Let $d=S_{i,j_{1}}-S_{i,j_{2}}$ denote the difference between two similarity scores in Eq. (5). The adaptive margin $m$ is then computed as follows:

### v2

@@ Abstract
Paired image-text data with subtle variations in-between (e.g., people holding surfboards vs. people holding shovels) hold the promise of producing Vision-Language Models with proper compositional understanding. Synthesizing such training data from generative models is a highly coveted prize due to the reduced cost of data collection. However, synthesizing training images for compositional learning presents three challenges: (1) efficiency in generating large quantities of images, (2) text alignment between the generated image and the caption in the exact place of the subtle change, and (3) image fidelity in ensuring sufficient similarity with the original real images in all other places. We propose SPARCL (Synthetic Perturbations for Advancing Robust Compositional Learning), which integrates image feature injection into a fast text-to-image generative model, followed by an image style transfer step, to meet the three challenges. Further, to cope with any residual issues of text alignment, we propose an adaptive margin loss to filter out potentially incorrect synthetic samples and focus the learning on informative hard samples. Evaluation on four compositional understanding benchmarks demonstrates that SPARCL significantly improves the compositionality of CLIP, boosting the average accuracy of the CLIP base model by over 8% across all benchmarks and outperforming state-of-the-art methods by 2% on three benchmarks.
@@ 1 Introduction
Current Vision-Language Models (VLMs) still face limitation in accurately interpreting compositional relationships between objects and attributes, as demonstrated by numerous evaluations [88, 113, 106, 62, 32]. This limitation primarily stems from the absence of subtle variations in the training data [42] (e.g., the subtle variations between the two captions in Figure 1 (a)). As a result, it becomes possible to maximize empirical image-caption alignment using shortcut features [23] rather than genuinely learning nuanced distinctions. While collecting training samples with subtle variations could enhance compositionality, this approach is time-consuming and labor-intensive, rendering it impractical at scale.
@@
However, generating and training on synthetic data presents two key challenges. The first is the difficulty of efficiently and accurately creating precise variations in synthetic images. To generate large-scale training data with precise variations, the image generation process must meet three criteria: efficiency in producing large quantities of images, text alignment of the generated image to the corresponding caption, which is subtly changed from the original caption, and image fidelity in ensuring the generated image otherwise closely matches their real counterpart. Nevertheless, current image generation methods struggle to meet all three criteria simultaneously. Per-sample optimization methods [44, 94, 112, 82] lack efficiency, while zero-shot image editing methods [63, 6, 22] and text-to-image (T2I) models [75, 78] often fail to achieve proper text alignment and image fidelity, respectively, as illustrated in Figure 1 (a). The inaccurate synthetic variations produced by these models can mislead the learning of VLMs. To alleviate this issue, we take a fast T2I model, which excels in efficiency and text alignment, and inject real image features into the text prompt features in order to enhance image fidelity. By combining this approach with AdaIN [35], we manage to substantially improve the image generation process per the three criteria.
Still, the generated data do not perfectly meet the text alignment criterion, and we propose to deal with the remaining problems with an innovative approach to model training. As Figure 1 (b) shows, the similarity between synthetic positive and negative pairs may vary, resulting in a mixture of hard text-image pairs, easy pairs, and incorrect pairs. Instead of treating all positive and negative samples uniformly, we propose a novel loss that differentiates between positive, hard negative, and easy negative samples. Further, we propose an adaptive margin that helps to filter out potentially incorrect generations and focus the learning on informative hard samples.
@@
In summary, we propose SPARCL, which integrates image feature injection into a fast T2I model to improve the quality of synthetic variations in images and employs an adaptive margin loss to leverage the varying alignment quality in synthetic samples for training VLMs. Evaluations on four compositional understanding benchmarks demonstrate that SPARCL significantly enhances VLM compositionality, improving CLIP by over 5% on VL-CheckList [113] and 7% on SugarCrepe [32], while surpassing state-of-the-art methods by 1% and 2% on the two benchmarks. The main contributions of this paper are as follows: (1) we propose image feature injection to enhance the quality of synthetic variations in images, which provide valuable training data to improve the compositionality of VLMs; (2) we introduce an adaptive margin loss to leverage varying levels of cross-modal alignment in synthetic samples to effectively differentiate positive and negative samples; (3) experimental results validate that SPARCL significantly improves the compositional understanding capabilities of CLIP models.
@@ 2.1 Limitations in Compositionality of VLMs
While VLMs excel in many multi-modal tasks [74, 40, 101, 84, 52, 53, 115], they still struggle with compositional understanding—the ability to interpret novel combinations of known visual and textual components. Benchmarks like What’sUp [42] reveal difficulties in understanding spatial relationships, SPEC [70] highlights issues with object size, position, and count, and ARO [106] uncovers limitations in understanding attributes, relations, and word order. Winoground [88], SNARE [96], and VL-CheckList [113] also expose these shortcomings. SugarCrepe [32] addresses hackable biases in prior benchmarks, where text-only models achieve artificially high performance, by introducing fluent and meaningful hard negatives. Their findings suggest that previous benchmarks overestimated compositional understanding. Building on this, SugarCrepe++ [19] further introduces semantically equivalent but lexically varied captions as hard positives and shows the difficulties of VLMs in distinguishing between lexical variations.
@@ 2.2 Improving Compositionality of VLMs
Prior approaches to improving the compositionality of VLMs can be broadly classified into the following categories: (1) Leveraging detailed image captions: Detailed captions from dense captioning models [17, 54], simulation platforms [9], and video annotations [45] are collected to train VLMs. However, these samples often lack pairs with subtle variations, limiting their contribution to compositionality. (2) Distilling from pretrained models: SDS-CLIP [3], SF-CLIP [80] and IL-CLIP [114] distill knowledge from pretrained image generation models and visual-language foundation models. IL-CLIP [114] refines representations through iterative learning with pretrained vision and language agents. However, pretrained models also face limitations in compositionality [34, 90]. (3) Incorporating structural knowledge: MosaiCLIP [85] and Structure-CLIP [36] incorporate scene graph knowledge in text features. CLIP-SGVL [28] and 3VL [103] train VLMs to predict scene graphs. [66] utilizes scene graphs as prompts to elicit compositional knowledge from VLMs without further training. However, these methods rely on models trained with expensive dense structure annotations (e.g., scene graphs). (4) Utilizing synthetic negative samples: Rule-based tools [30] or large language models (LLMs) [15, 73] are used to generate negative captions by editing real captions [106, 18, 11, 86, 83, 68, 109, 10]. Image generation models [75] are also used to create or edit images for training [79, 71, 108, 49]. Despite their utility, efficiently generating large amount of images with precise variations remains challenging due to the inherent limitations of generative models. (5) Applying fine-grained alignment constraints: MCD [46] enforces multi-scale alignment across images with varying augmentations and the corresponding text captions. SPARC [5] learns local alignment by associating each text token with a group of local image patches. CE-CLIP [109] applies intra-modal contrastive loss and cross-modal ranking loss to improve alignment. However, uniform supervision applied to synthetic samples with varying alignment quality limits their effectiveness. In this paper, we propose SPARCL to address two key challenges in learning from synthetic data: the difficulty in accurately creating precise variations and the inconsistency in cross-modal alignment quality in synthetic data.
@@ 2.3 Training with Synthetic Data
Synthetic data for training machine learning models have been studied in various fields [67, 76, 77, 48, 99, 26, 64, 14, 93]. Synthetic data generated through simulations and graphics engines supports a wide range of tasks [16, 72, 95]. However, these synthetic datasets often diverge significantly from real-world data. Recent advances in generative models [7, 110, 91, 43, 12, 75, 78, 105, 61] have made it possible to synthesize data that more closely resembles real-world scenarios. Synthetic data are widely used in both language tasks [21, 31, 64, 65, 97, 98, 116, 87, 100, 60] and vision tasks [4, 25, 1, 81, 29, 92, 111, 39, 57, 2, 89]. Synthetic data are often noisy, necessitating noise-resistant training methods [41, 50, 13, 102, 51, 37], especially those designed for contrastive learning [56, 38]. In this paper, we generate multimodal samples with subtle variations and filter out potentially incorrect ones during training to improve the compositionality of VLMs.
Figure 2: An overview of SPARCL. (a) Starting with a real image-caption pair, we generate synthetic positive and negative pairs with subtle variations using an LLM and a fast T2I model. To improve the quality of subtle variations in synthetic images, we introduce image feature injection to reduce unintended variations from a standard T2I model (see Sec. 3.2). (b) We train the VLM using both real and synthetic samples. In addition to a sigmoid loss for distinguishing positive and negative pairs, we apply an adaptive margin loss that leverages varying alignment levels across training samples to learn informative nuanced distinctions (see Sec. 3.3).
3 SPARCL
To improve the compositional understanding abilities of VLMs, we propose Synthetic Perturbations for Advancing Robust Compositional Learning, or SPARCL, which generates multimodal samples with subtle variations from real samples and trains VLMs to learn nuanced differences through synthetic data. In the generation phase, SPARCL creates positive and negative captions with slight variations from real captions using an LLM, then generates images based on the real image and these modified captions through a fast T2I model. To enhance the quality of subtle variations in synthetic images, we introduce image feature injection, which integrates real image features into the text prompt features of the T2I model to improve fidelity to the real image, as detailed in Sec. 3.2. In the training phase, SPARCL employs an adaptive margin loss that leverages varying levels of multimodal alignment in the synthetic samples to effectively learn informative nuanced distinctions, as described in Sec. 3.3. The framework of SPARCL is shown in Figure 2.
@@ 3.1 Generating Negative and Positive Captions
Captions with subtle variations are crucial for learning compositional knowledge, as shown by previous work that generates captions by randomly swapping or replacing nouns and adjectives [106, 109]. However, manually designed generation rules often introduce nonsensical or grammatically incorrect artifacts, creating shortcut features [23] that obstruct true compositional understanding [32]. To address this issue, we use an LLM to generate natural synthetic captions. To further mitigate the impact of generative artifacts, we generate both negative and positive captions, ensuring that VLMs cannot easily differentiate negative captions from positive ones based solely on artifacts. We denote the $i^{th}$ real image-caption pair in the training set as $(\bm{I}_{i}^{r},\bm{T}_{i}^{r})$. Given a real caption $\bm{T}_{i}^{r}$, we prompt the LLM to generate a synthetic negative caption $\bm{T}_{i}^{sn}$ and a synthetic positive caption $\bm{T}_{i}^{sp}$, as specified by the prompts in Figure A1 in the Appendix.
@@ 3.2 Generating Images via Image Feature Injection
We extract features for a real image $\bm{I}_{i}^{r}$ and the corresponding synthetic caption $\bm{T}_{i}^{s}$ ($s\in\{sn,sp\}$) using two aligned feature encoders (e.g., CLIP image and text encoders). The image embedding $\bm{f}_{i}^{r}$ is the encoder output at the position of the [CLS] token. The text embedding is the text encoder output $\bm{e}_{i}^{s}=\langle\bm{e}_{i,j}^{s}\rangle_{j=1}^{L}$, where $L$ is the maximum sequence length. Let the index of the [EOS] token in $\bm{e}_{i}^{s}$ be $k_{i}^{s}$, $k_{i}^{s}\leq L$. That is, the padding token [PAD] is used as input to the text encoder at positions between $k_{i}^{s}$ and $L$. The key step of SPARCL is to replace the text embeddings at those positions with the image embedding, producing a new sequence of text embeddings, $\hat{\bm{e}}_{i}^{s}$,
$\hat{\bm{e}}_{i}^{s}=\langle\bm{e}_{i,1}^{s},\bm{e}_{i,2}^{s},\ldots,\bm{e}_{i,k_{i}}^{s},\underbrace{\bm{f}_{i}^{r},\ldots,\bm{f}_{i}^{r}}_{L-k_{i}^{s}\text{ times}}\rangle.$ (1)
The embeddings $\hat{\bm{e}}_{i}^{s}$ is then used as input to the T2I model to generate a synthetic image $\tilde{\bm{I}}_{i}^{s}$. This process, depicted in Figure 2, reduces unintended variations in synthetic images generated by the standard T2I model and enhances image fidelity to real images, thereby improving the quality of synthetic subtle variations.
Style Transfer. To further reduce the domain gap between synthetic and real images, we use AdaIN [35] to transfer the style of the synthetic image to that of the real image. We use a pretrained AdaIN encoder to extract content features from $\tilde{\bm{I}}_{i}^{s}$ and style features from $\bm{I}_{i}^{r}$. We normalize the content features using instance normalization and then scale and shift them with the mean and variance of the style features. The transformed features are fed into a pretrained AdaIN decoder to generate $\bm{I}_{i}^{s}$, which is subsequently used for model training.
@@ 3.3 Training with Real and Synthetic Samples
After generating captions and images, each real image-caption pair $(\bm{I}_{i}^{r},\bm{T}_{i}^{r})$ is extended to $(\bm{I}_{i}^{r},\bm{T}_{i}^{r},\bm{I}_{i}^{sn},\bm{T}_{i}^{sn},\bm{I}_{i}^{sp},\bm{T}_{i}^{sp})$ by adding one synthetic negative and one synthetic positive pair. We then train the VLM using these extended samples. Given a batch of $n$ sample groups, $\{(\bm{I}_{i}^{r},\bm{T}_{i}^{r},\bm{I}_{i}^{sn},\bm{T}_{i}^{sn},\bm{I}_{i}^{sp},\bm{T}_{i}^{sp})\}_{i=1}^{n}$, we organize the images by concatenating all real, synthetic negative, and synthetic positive images into an image batch $\bm{I}^{B}$ of $3n$ images, and similar all captions into a caption batch $\bm{T}^{B}$ of $3n$ captions.
We calculate the similarity between each image-caption pair in the batch. The similarity between the $i^{\text{th}}$ image $\bm{I}_{i}^{B}$ and the $j^{\text{th}}$ caption $\bm{T}_{j}^{B}$ is given by $S_{i,j}=c(\mathcal{E}_{I}(\bm{I}_{i}^{B}),\mathcal{E}_{T}(\bm{T}_{j}^{B}))$, where $c(\cdot,\cdot)$ denotes cosine similarity, and $\mathcal{E}_{I}$ and $\mathcal{E}_{T}$ represent the image and text encoders of the VLM, respectively. We then define the ground-truth alignment variable $M_{i,j}$, which takes value $1$ if the image matches with the caption, and $-1$ otherwise. Clearly, the only positive image-caption pairs are $(\bm{I}_{i}^{r},\bm{T}_{i}^{r}),(\bm{I}_{i}^{sn},\bm{T}_{i}^{sn}),(\bm{I}_{i}^{sp},\bm{T}_{i}^{sp}),(\bm{I}_{i}^{r},\bm{T}_{i}^{sp})$ and $(\bm{I}_{i}^{sp},\bm{T}_{i}^{r})$. To account for multiple positive associations of an image or a caption, we apply a sigmoid-based contrastive loss [107, 8] to encourage higher similarity for positive pairs and lower similarity for negative pairs,
@@
Adaptive Margin Loss. Despite best efforts toward controlling generative models, synthetic samples may still have variable quality in terms of the alignment between text and imagery. That is, purported positive (resp. negative) pairs may not be semantically similar (resp. dissimilar). To account for variation in synthetic data quality, we propose to differentiate between real data and synthetic data and between easy negatives and hard negatives. Further, we propose an adaptive margin that filters out potential incorrect samples and prioritizes learning from hard samples.
For each image in a batch, we define four sets of captions, a positive set $\mathbb{P}$, a hard negative set $\mathbb{N}_{h}$, an easy negative set $\mathbb{N}_{e}$, and a real negative set $\mathbb{N}_{r}$, which represent different levels of alignment within the batch. If the image is a real or synthetic positive image,
  • $\mathbb{P}$ contains the real positive captions $\bm{T}_{i}^{r}$ and the synthetic positive captions $\bm{T}_{i}^{sp}$ of the current, $i$-th image.
  • $\mathbb{N}_{h}$ contains the synthetic negative captions $\bm{T}_{i}^{sn}$ of the current image.
  • $\mathbb{N}_{e}=\{\bm{T}_{j}^{r},\bm{T}_{j}^{sp},\bm{T}_{j}^{sn}|i\neq j\}$ contains all real and synthetic captions from all other images.
  • $\mathbb{N}_{r}=\{\bm{T}_{j}^{r}|i\neq j\}$ contains all real captions belonging to the other images.
If the image is a synthetic negative image, the sets $\mathbb{N}_{e}$ and $\mathbb{N}_{r}$ are unchanged but $\mathbb{P}$ and $\mathbb{N}_{h}$ differ:
  • $\mathbb{P}$ contains the synthetic negative captions $\bm{T}_{i}^{sn}$.
  • $\mathbb{N}_{h}$ contains $\bm{T}_{i}^{r}$ and $\bm{T}_{i}^{sp}$.
We define the margin loss for the $i^{th}$ image as:
$\begin{split}L_{mar,i}^{I}&=\frac{1}{\lvert\mathbb{P}\rvert\cdot\lvert\mathbb{N}_{h}\rvert}\sum_{j_{1}\in\mathbb{P},j_{2}\in\mathbb{N}_{h}}\text{max}(0,m+S_{i,j_{2}}-S_{i,j_{1}})\\ &+\frac{1}{\lvert\mathbb{N}_{h}\rvert\cdot\lvert\mathbb{N}_{e}\rvert}\sum_{j_{1}\in\mathbb{N}_{h},j_{2}\in\mathbb{N}_{e}}\text{max}(0,m+S_{i,j_{2}}-S_{i,j_{1}})\\ &+\frac{\alpha}{\lvert\mathbb{P}\rvert\cdot\lvert\mathbb{N}_{r}\rvert}\sum_{j_{1}\in\mathbb{P},j_{2}\in\mathbb{N}_{r}}\text{max}(0,m+S_{i,j_{2}}-S_{i,j_{1}}).\end{split}$ (3)
This loss encourages positive pairs to have higher similarity scores than hard negative pairs, which should have higher similarity than easy negative pairs. Additionally, a weight $\alpha>1$ is applied to comparisons between positive pairs and real negative pairs to emphasize these comparisons, as real samples are generally correct, making comparisons involving them more reliable.
Lastly, we propose an adaptive margin in the above loss. Let $d=S_{i,j_{1}}-S_{i,j_{2}}$ denote the difference between two similarity scores in Eq. (3). The adaptive margin $m$ is then computed as follows: